OpenStudy (anonymous):

Trying to solve the equation a^2(d^2u/dx^2) = (d^2u/dt^2) with the given conditions: x>0, t>0, u(0,t) = f(t), u(x,0) = 0, u-subt(x,0) = 0 and the lim as x-->infinity of u(x,t) = 0 for t>0 Need to use laplace...ideas?

OpenStudy (amistre64):

... my idea would be to use Laplace then :)

OpenStudy (amistre64):

trying to read thru the notation tho is confusing me

OpenStudy (amistre64):

\[L\{a^2\frac{d^2u}{dx^2}=\frac{d^2u}{dt^2}\}\]

OpenStudy (amistre64):

is a^2 consider a constant?

OpenStudy (amistre64):

\[L\{a^2\frac{d^2u}{dx^2}\}=s^2U(s)-su(0)-u'(0)\] i might have got the u and u' mixed up